OpenStudy (anonymous):

Prove convergence or divergence for the series: ∑∞ n=0 [(2n+1)/3n+2)] ∑∞ n=0 [(2n-1)/(3n3 + 2n + 4)] ∑∞ n=1 [absolute value of (sin n)] / n2

OpenStudy (anonymous):

State test used as well plz

OpenStudy (anonymous):

For the first one, use the test for divergence. For a sum to converge we must have that: \[\lim_{k \rightarrow \infty}a_k=0\] For this we have: \[\lim_{n \rightarrow \infty}\frac{2n+1}{3n+2}=2/3 \ne 0 \implies \sum_j \frac{2j+1}{3n+2} D.N.C.\] Does not converge.

OpenStudy (anonymous):

Thanks can you do this for the others as well?

OpenStudy (anonymous):

Well for the second one you want to see what it does in the long term, as n gets large it is dominated by: \[\frac{2}{3n^2}\] So simply pick a CONVERGENT series that is BIGGER. Say... \[\sum_k \frac{1000}{k^2}\] Which is GREATER than your series for large n. Therefore by limit comparison SINCE 1000/k^2 CONVERGES, then your series converges. http://www.wolframalpha.com/input/?i=sum+n%3D0+to+infinity+of+(2n-1)%2F(3n%5E3%2B2n%2B4)

OpenStudy (anonymous):

Thanks so much and for the third one?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{|\sin(n)|}{n^2}\] Well we know that sin(n) goes between -1 and 1 and the absolute value restricts that to just between 0 and 1, this means that a series such as 3/n^2 would be bigger therefore: \[\sum_n \frac{|\sin(n)|}{n^2} < \sum_n \frac{3}{n^2} \] and 3/n^2 converges so by limit comparison the other converges. http://www.wolframalpha.com/input/?i=sum+from+n+%3D+1+to+infinity+of+%7Csin(n)%7C%2Fn%5E2 It apparently converges to like 1.3ish

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

No problem :P